# Reinforcement Learning in Vision: A Survey

### v2

@@ 3 Reinforcement Learning in Vision
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, rounded corners, align=center, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8.8em,font=,, where level=2text width=10.2em,font=,, where level=3text width=7.2em,font=,, where level=4text width=35.5em,font=,, [ Reinforcement Learning in Vision, ver, color=carminepink!100, fill=carminepink!15, text=black [ Multimodal Large
Language Models (§3.1), color=lighttealblue!100, fill=lighttealblue!100, text=black [ Conventional RL-based
Frameworks(§3.1.1), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  RePIC (Oh et al., 2025), GoalLadder (Zakharov & Whiteson, 2025), Drive-R1 (Li et al., 2025e),
VLM-R1 (Shen et al., 2025b), GRPO-CARE (Chen et al., 2025b), Q-Ponder (Cai et al., 2025c),
MoDoMoDo (Liang et al., 2025), V-Triune (Ma et al., 2025),Vision-R1 (Zhan et al., 2025b),
ProxyThinker (Xiao et al., 2025b), Jigsaw-R1 (Wang et al., 2025k), SRPO (Wan et al., 2025),
R1-Onevision (Yang et al., 2025f), Reason-RFT (Tan et al., 2025a), Segagent (Zhu et al., 2025c),
VisualQuality-R1 (Wu et al., 2025c), Zhang et al. (Zhang et al., 2025a), SATORI-R1 (Shen et al., 2025a),
Skywork r1v2 (Wang et al., 2025e), Yang et al. (Yang et al., 2025a), etc. , cause_leaf, text width=38em ] ] [ Spatial & 3D Perception
3.1.2), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  Omni-R1 (Zhong et al., 2025), DIP-R1 (Park et al., 2025b), Metaspatial (Pan & Liu, 2025),
Bindgpt (Zholus et al., 2025), Scene-R1 (Yuan et al., 2025c), Perception-R1 (Yu et al., 2025),
VisRL (Chen et al., 2025h), ViCrit (Wang et al., 2025g), VisionReasoner (Liu et al., 2025e),
ViGoRL (Sarch et al., 2025), Visual-RFT (Liu et al., 2025f), etc. , cause_leaf, text width=38em ] ] [ Image Reasoning (§3.1.3), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Think about Image, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g.,  SVQA-R1 (Wang & Ling, 2025), VL-GenRM (Zhang et al., 2025d), RACRO (Gou et al., 2025)
, EasyARC (Unsal & Akkus, 2025), STAR-R1 (Li et al., 2025g), Visionary-R1 (Xia et al., 2025),
UniVG-R1 (Bai et al., 2025), EchoInk-R1 (Xing et al., 2025), WeThink (Yang et al., 2025b),
G1 (Chen et al., 2025a), GThinker (Zhan et al., 2025a), Observe-R1 (Guo et al., 2025c),
Mm-eureka (Meng et al., 2025), Perception-R1 (Xiao et al., 2025a), MiMo-VL (Xiaomi, 2025),
SearchExpert (Li et al., 2025b), etc. , cause_leaf, text width=38em ] ] [ Think with Image, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g.,  Visual Planning (Xu et al., 2025b), GRIT (Fan et al., 2025), VILASR (Wu et al., 2025a),
BRPO (Chu et al., 2025), Ground-R1 (Cao et al., 2025), Pixel Reasoner (Su et al., 2025a),
DeepEyes (Zheng et al., 2025), TACO (Kan et al., 2025), VRAG-RL (Wang et al., 2025f),
VisTA (Huang et al., 2025),TGI (Chern et al., 2025), Chain-of-Focus (Zhang et al., 2025g),
Openthinkimg (Su et al., 2025b), Active-O3 (Zhu et al., 2025d), etc. , cause_leaf, text width=38em ] ] ] [ Video Reasoning(§3.1.4), color=lighttealblue!100, fill=lighttealblue!60, text=black [ e.g.,  VQ-Insight (Zhang et al., 2025h), TW-GRPO (Dang et al., 2025), EgoVLM (Vinod et al., 2025),
VAU-R1 (Zhu et al., 2025b), DeepVideo-R (Park et al., 2025a), TimeMaster (Zhang et al., 2025e),
VideoR1 (Feng et al., 2025), VideoChat-R1 (Li et al., 2025d), Temporal-RLT (Li et al., 2025a), etc. , cause_leaf, text width=38em ] ] ] [ Visual Generation(§3.2), color=lightplum!100, fill=lightplum!100, text=black [ Image Generation(§3.2.1), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  ImageReward (Xu et al., 2023), ReasonGen-R1 (Zhang et al., 2025i), FocusDiff (Pan et al., 2025),
Dpok (Fan et al., 2023a), LOOP (Gupta et al., 2025), Prefpaint (Liu et al., 2024a),
RLRF (Rodriguez et al., 2025b), GoT-R1 (Duan et al., 2025), D-Fusion (Hu et al., 2025b),
Black et al., (Black et al., 2023), DiffusionDPO (Wallace et al., 2024), Diffusion-KTO (Li et al., 2024b),
DiffusionRPO (Gu et al., 2024), Miao et al., (Miao et al., 2024b), Parrot (Lee et al., 2024),
DPG-T2I (Wei et al., 2024), RPO (Miao et al., 2024a), RePrompt (Wu et al., 2025b),
RLCM (Oertell et al., 2024), B2-DiffuRL (Hu et al., 2025a), DiffPPO (Xiao et al., 2024),
Simplear(Wang et al., 2025c), LLaVA-Reward (Zhou et al., 2025b), Flow-grpo (Liu et al., 2025a),
Gallici et al. (Gallici & Borde, 2025), TexForce (Chen et al., 2024a), etc. , detect_leaf, text width=38em ] ] [ Video Generation(§3.2.2), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  DanceGRPO (Xue et al., 2025), InfLVG (Fang et al., 2025), Phys-AR (Lin et al., 2025),
VideoReward (Liu et al., 2025b), TeViR (Chen et al., 2025d), GAPO (Zhu et al., 2025a),
Instructvideo (Yuan et al., 2024), Videoscore (He et al., 2024b),
Gradeo (Mou et al., 2025), VideoRM (Wu et al., 2024), etc. , detect_leaf, text width=38em ] ] [ 3D Generation(§3.2.3), color=lightplum!100, fill=lightplum!60, text=black [ e.g.,  DreamCS (Zou et al., 2025), Dreamreward (Ye et al., 2024), DreamDPO (Zhou et al., 2025c),
Nabla-R2D3 (Liu et al., 2025d), etc. , detect_leaf, text width=38em ] ] ] [ Unified Model(§3.3), color=harvestgold!100, fill=harvestgold!100, text=black [ Unified RL(§3.3.1), color=harvestgold!100, fill=harvestgold!60, text=black [ e.g.,  UniRL (Mao et al., 2025), CoRL (Jiang et al., 2025b), SelfTok (Wang et al., 2025a),
Hermesflow (Yang et al., 2025c), UnifiedReward (Wang et al., 2025j),
UnifiedReward-Think (Wang et al., 2025i), etc. , detect_harvestgold, text width=38em ] ] [ Task-specific RL(§3.3.2), color=harvestgold!100, fill=harvestgold!60, text=black [ e.g.,  Vargpt-v1. 1 (Zhuang et al., 2025), Emu3 (Wang et al., 2024b), X-Omni (Geng et al., 2025), etc. , detect_harvestgold, text width=38em ] ] ] [ Vision Language Action
Models(§3.4), color=burlywood!100, fill=burlywood!100, text=black [ GUI Interaction(§3.4.1), color=burlywood!100, fill=burlywood!60, text=black [ e.g.,  GUI-R1 (Luo et al., 2025), SE-GUI (Yuan et al., 2025b), UI-R1 (Lu et al., 2025),
UIShift (Gao et al., 2025b), AgentCPM-GUI (Zhang et al., 2025j), MobileGUI-RL (Shi et al., 2025),
ProgRM (Zhang et al., 2025b), Mobile-R1 (Gu et al., 2025), GTA1 (Yang et al., 2025e),
LPO (Tang et al., 2025), Ui-tars (Qin et al., 2025), RUIG (Zhang et al., 2023),
Appvlm (Papoudakis et al., 2025) etc. , detect_burlywood, text width=38em ] ] [ Visual Navigation(§3.4.2), color=burlywood!100, fill=burlywood!60, text=black [ e.g.,  OctoNav-R1 (Gao et al., 2025a), More (Zhao et al., 2025), RAPID (Kim et al., 2025),
VLN-R1 (Qi et al., 2025), Flare (Hu et al., 2024a), etc. , detect_burlywood, text width=38em ] ] [ Visual Manipulation(§3.4.3), color=burlywood!100, fill=burlywood!60, text=black [ e.g.,  TGRPO (Chen et al., 2025g), RFTF (Shu et al., 2025), RLVLA (Liu et al., 2025c),
VLA-RL (Lu et al., 2025b), ConRFT (Chen et al., 2025e), iRe-VLA (Guo et al., 2025b),
RIPT-VLA (Tan et al., 2025b), ReinBot (Zhang et al., 2025c),
Flare (Hu et al., 2024a), etc. , detect_burlywood, text width=38em ] ] ] ]
@@ 3.1.1 Conventional RL-based MLLMs
We refer to conventional RL-based MLLMs as approaches that apply reinforcement learning primarily to align a vision–language backbone with verifiable, task-level rewards, without explicitly modeling multi-step chain-of-thought reasoning. Typical works RePIC (Oh et al., 2025), GoalLadder (Zakharov & Whiteson, 2025), Drive-R1 (Li et al., 2025e) and VLM-R1 (Shen et al., 2025b) replace preference models with deterministic validators (e.g., exact-match, IoU, BLEU) and optimize the policy by GRPO/PPO variants under a KL regulariser. This design yields stable value-free training, improves zero-shot robustness on captioning, grounding and autonomous-driving benchmarks, and substantially reduces the annotation cost typically incurred by supervised fine-tuning.
Recent extensions demonstrate the flexibility of this paradigm. GRPO-CARE (Chen et al., 2025b) introduces consistency-aware group normalization to mitigate reward variance, while Q-Ponder (Cai et al., 2025c) adds a pondering controller. From a data perspective, MoDoMoDo formulates a multi-domain mixture optimization that predicts reward distributions and selects optimal curricula (Liang et al., 2025); V-Triune further unifies perception and reasoning tasks within a single triple-objective pipeline, empirically validating that rule-based RL scales to diverse visual signals (Ma et al., 2025). Collectively, these studies indicate that (i) verifiable rewards can serve as a low-cost alternative to human feedback, (ii) group-relative objectives offer higher training stability than token-level PPO on heterogeneous visual tasks, and (iii) curriculum or data-mixture scheduling is emerging as a key ingredient for broad generalization.
@@ 3.1.2 Spatial and 3D Perception
2D perception. Perception centric works applies RL to sharpen object detection, segmentation and grounding without engaging in lengthy chain–of–thought reasoning. Omni-R1 (Zhong et al., 2025) introduces a two-system (global–local) GRPO pipeline that verifies predictions via rule-based metrics, yielding notable improvements on region-level benchmarks and emotion-recognition tasks. DIP-R1 (Park et al., 2025b) further decomposes perception into step-wise “inspect $\rightarrow\!$ observe $\rightarrow\!$ act” cycles, where each stage receives deterministic IoU or counting-based rewards to boost fine-grained detection. Perception-R1 (Yu et al., 2025) revisits the effect of GRPO on a spectrum of detection and OCR datasets. Complementing these, VisRL (Chen et al., 2025h) frames intention-guided focus selection as an RL sub-policy, eliminating expensive region labels and consistently outperforming supervised strong baselines on visual grounding tasks.
3D perception. Beyond 2-D, several studies leverage RL to align multimodal models with physically consistent 3-D layouts. MetaSpatial (Pan & Liu, 2025) employs rendered depth/IoU rewards to refine spatial reasoning for AR/VR scene generation, whereas Scene-R1 (Yuan et al., 2025c) couples video-grounded snippet selection with a two-stage grounding policy to learn 3-D scene structure without point-level supervision. At molecular scale, BindGPT (Zholus et al., 2025) treats atom placement as sequential actions and uses binding-affinity estimators as verifiable rewards, demonstrating the scalability of perception-focused RL to 3-D biochemical design. Collectively, these approaches underscore a common recipe: (i) formulate detection/segmentation/3-D alignment as Markov decision problems, (ii) craft deterministic spatial rewards (e.g., IoU, depth consistency, binding energy), and (iii) fine-tune pretrained VLM backbones via GRPO/PPO for stable perception enhancement—thereby differentiating themselves from reasoning-oriented RL variants.
@@ 3.1.3 Image Reasoning
Think about Image. Early think about image works for spatial VQA employs view-consistent or transformation-invariant objectives, such as SVQA-R1 (Wang & Ling, 2025) and STAR-R1 (Li et al., 2025g). VL-GenRM (Zhang et al., 2025d) and RACRO (Gou et al., 2025) refine preference data or caption rewards to curb hallucinations. Benchmark-oriented efforts such as EasyARC (Unsal & Akkus, 2025) offer procedurally generated, fully verifiable tasks that suit outcome-based reinforcement learning. To mitigate shortcut reliance and improve generality, Visionary-R1 (Xia et al., 2025) enforces image interpretation before reasoning, whereas UniVG-R1 (Bai et al., 2025) unifies referring, captioning, and detection by coupling a grounding corpus with GRPO fine-tuning. Extensions such as EchoInk-R1 (Xing et al., 2025) further enrich visual reasoning by integrating audio–visual synchrony under GRPO optimization. Meanwhile, curriculum-driven frameworks—WeThink, G1, GThinker, and Observe-R1 progressively increase task complexity or introduce re-thinking cues (e.g., difficulty ladders, multimodal format constraints) to cultivate deeper and more structured reasoning capabilities in MLLMs. These methods show that language only RL with well-designed visual correctness rewards significantly improves model accuracy, robustness, and out-of-distribution performance.
Think with Image. Early think with image works grounds reasoning via discrete region–level operations: GRIT (Fan et al., 2025) interleaves bounding-box tokens with language and trains under GRPO to maximize both answer correctness and box fidelity, while VILASR (Wu et al., 2025a) generalizes this idea to multi-view and video settings, enforcing cross-view spatial consistency. Ground-R1 (Cao et al., 2025) and BRPO (Chu et al., 2025) adopt two-stage pipelines that first highlight evidence regions (via IoU-based or reflection rewards) before verbal reasoning. A complementary thread explores pixel-space or sequence-level manipulation. Visual Planning (Xu et al., 2025b) replaces text chains with imagined image roll-outs rewarded by downstream task success; Pixel Reasoner (Su et al., 2025a) augments the action space with crop, erase and paint primitives and balances exploration through curiosity-driven rewards, whereas DeepEyes (Zheng et al., 2025) shows that end-to-end RL can spontaneously induce such visual thinking behaviours. Finally, TACO (Kan et al., 2025) introduces a think–answer consistency objective that resamples long visual–verbal chains until their intermediate edits align with the final answer. Together, these systems demonstrate that explicitly generating or editing visual artefacts during reasoning optimized via GRPO or R1-style outcome RL—yields more faithful, interpretable and robust image understanding than language only approaches.

### v3

@@ 3 Reinforcement Learning in Vision
[…]
Reinforcement Learning in Vision
Multimodal Large Language Models (§3.1) Conventional RL-based Frameworks(§3.1.1) e.g.,  RePIC [oh2025repic], GoalLadder [zakharov2025goalladder], Drive-R1 [li2025drive], VLM-R1 [shen2025vlm], GRPO-CARE [chen2025grpo], Q-Ponder [cai2025q], MoDoMoDo [liang2025modomodo], V-Triune [ma2025one],Vision-R1 [zhan2025vision], ProxyThinker [xiao2025proxythinker], Jigsaw-R1 [wang2025jigsaw], SRPO [wan2025srpo], R1-Onevision [yang2025r1], Reason-RFT [tan2025reason], Segagent [zhu2025segagent], VisualQuality-R1 [wu2025visualquality], Zhang et al. [zhang2025improving], SATORI-R1 [shen2025satori], Skywork r1v2 [wang2025skywork], Yang et al. [yang2025self], THOR [chang2025thor] Caprl [xing2025caprl], LENS [zhu2025lens], Docr1 [xiong2025docr1], Bigcharts-r1 [masry2025bigcharts], etc. Spatial & 3D Perception 3.1.2) e.g.,  Omni-R1 [zhong2025omni], DIP-R1 [park2025dip], Metaspatial [pan2025metaspatial], Bindgpt [zholus2025bindgpt], Scene-R1 [yuan2025scene], Perception-R1 [yu2025perception], VisRL [chen2025visrl], ViCrit [wang2025vicrit], VisionReasoner [liu2025visionreasoner], ViGoRL [sarch2025grounded], Visual-RFT [liu2025visual], etc. Image Reasoning (§3.1.3) Think about Image e.g.,  SVQA-R1 [wang2025svqa], VL-GenRM [zhang2025vl], RACRO [gou2025perceptual], EasyARC [unsal2025easyarc], STAR-R1 [li2025star], Visionary-R1 [xia2025visionary], UniVG-R1 [bai2025univg] , EchoInk-R1 [xing2025echoink], WeThink [yang2025wethink], G1 [chen2025g1], GThinker [zhan2025gthinker], Observe-R1 [guo2025observe], Mm-eureka [meng2025mm], Perception-R1 [xiao2025advancing], MiMo-VL [coreteam2025mimovltechnicalreport], SearchExpert [li2025enhancing], etc. Think with Image e.g.,  Visual Planning [xu2025visual], GRIT [fan2025grit], VILASR [wu2025reinforcing], BRPO [chu2025qwen], Ground-R1 [cao2025ground], Pixel Reasoner [su2025pixel], DeepEyes [zheng2025deepeyes] , TACO [kan2025taco], VRAG-RL [wang2025vrag], VisTA [huang2025visualtoolagent],TGI [chern2025thinking], Chain-of-Focus [zhang2025chain], Openthinkimg [su2025openthinkimg], Active-O3 [zhu2025active] , RRVF [chen2025learning], Visionthink [yang2025visionthink], etc. Video Reasoning(§3.1.4) e.g.,  VQ-Insight [zhang2025vq], TW-GRPO [dang2025reinforcing], EgoVLM [vinod2025egovlm], VAU-R1 [zhu2025vau], DeepVideo-R [park2025deepvideo], TimeMaster [zhang2025timemaster], VideoR1 [videor1], VideoChat-R1 [li2025videochat], Temporal-RLT [li2025reinforcement], Moss-chatv [tao2025moss], Tempsamp-r1 [li2025tempsamp], Video-mtr [xie2025video], Thinking with videos [zhang2025thinking], MIRO [dufour2025miro], Chunk-GRPO [luo2025sample], Onereward [gong2025onereward], Uso [wu2025uso], Pref-grpo [wang2025pref] , Tempflow-grpo [he2025tempflow], etc. Visual Generation(§3.2) Image Generation(§3.2.1) e.g.,  ImageReward [xu2023imagereward], ReasonGen-R1 [zhang2025reasongen], FocusDiff [pan2025focusdiff], Dpok [fan2023dpok], LOOP [gupta2025simple], Prefpaint [liu2024prefpaint], RLRF [rodriguez2025rendering] , GoT-R1 [duan2025got], D-Fusion [hu2025d], Black et al., [black2023ddpo], DiffusionDPO [wallace2024diffusion], Diffusion-KTO [li2024aligning],DiffusionRPO [gu2024diffusion], Miao et al., [miao2024training], Parrot [lee2024parrot], DPG-T2I [wei2024powerful], RPO [miao2024subject], RePrompt [wu2025reprompt], RLCM [oertell2024rl], B2-DiffuRL [hu2025towards], DiffPPO [xiao2024diffppo], Simplear[wang2025simplear], LLaVA-Reward [zhou2025multimodalllmscustomizedreward], Flow-grpo [liu2025flow], Gallici et al. [gallici2025fine], TexForce [chen2024enhancing], etc. Video Generation(§3.2.2) e.g.,  DanceGRPO [xue2025dancegrpo], InfLVG [fang2025inflvg], Phys-AR [lin2025reasoning], VideoReward [liu2025improving], TeViR [chen2025tevir], GAPO [zhu2025aligning], Instructvideo [yuan2024instructvideo], Videoscore [he2024videoscore], Gradeo [mou2025gradeo], VideoRM [wu2024boosting], etc. 3D Generation(§3.2.3) e.g.,  DreamCS [zou2025dreamcs], Dreamreward [ye2024dreamreward], DreamDPO [zhou2025dreamdpo], Nabla-R2D3 [liu2025nabla], etc. Unified Model(§3.3) Unified RL(§3.3.1) e.g.,  UniRL [mao2025unirl], CoRL [jiang2025co], SelfTok [wang2025discrete], Hermesflow [yang2025hermesflow], UnifiedReward [wang2025unified], UnifiedReward-Think [wang2025unified1], etc. Task-specific RL(§3.3.2) e.g.,  Vargpt-v1. 1 [zhuang2025vargpt], Emu3 [wang2024emu3], X-Omni [geng2025x], etc. Vision Language Action Models(§3.4) GUI Interaction(§3.4.1) e.g.,  GUI-R1 [luo2025gui], SE-GUI [yuan2025enhancing], UI-R1 [lu2025ui], UIShift [gao2025uishift], AgentCPM-GUI [zhang2025agentcpm], MobileGUI-RL [shi2025mobilegui], ProgRM [zhang2025progrm] , Mobile-R1 [gu2025mobile], GTA1 [yang2025gta1], LPO [tang2025lpo], Ui-tars [qin2025ui], RUIG [zhang2023reinforced], Appvlm [papoudakis2025appvlm], DigiRL [bai2024digirl], ARPO [arpo], etc. Visual Navigation(§3.4.2) e.g.,  OctoNav-R1 [gao2025octonav], More [zhao2025more], RAPID [kim2025rapid], VLN-R1 [qi2025vln], Flare [hu2024flare], IRL-VLA [jiang2025irl], Embodied-R [zhao2025embodied], EmbCLIP-Agent [eftekhar2023selective], etc. Visual Manipulation(§3.4.3) e.g.,  TGRPO [chen2025tgrpo], RFTF [shu2025rftf], RLVLA [liu2025can], VLA-RL [lu2025vla], ConRFT [chen2025conrft], iRe-VLA [guo2025improving], RIPT-VLA [tan2025interactive], ReinBot [zhang2025reinbot], VIKI-R [kang2025viki], Robot-R1 [kim2025robot], RLDG [xu2024rldg], HIL-SERL [luo2024precise], Modem [hansen2022modem], VPG [zeng2018learning], etc.
@@ 3.1.1 Conventional RL-based MLLMs
We refer to conventional RL-based MLLMs as approaches that apply reinforcement learning primarily to align a vision–language backbone with verifiable, task-level rewards, without explicitly modeling multi-step chain-of-thought reasoning. Typical works RePIC [oh2025repic], GoalLadder [zakharov2025goalladder], Drive-R1 [li2025drive] and VLM-R1 [shen2025vlm] replace preference models with deterministic validators (e.g., exact-match, IoU, BLEU) and optimize the policy by GRPO/PPO variants under a KL regulariser. This design yields stable value-free training, improves zero-shot robustness on captioning, grounding and autonomous-driving benchmarks, and substantially reduces the annotation cost typically incurred by supervised fine-tuning.
Recent extensions demonstrate the flexibility of this paradigm. GRPO-CARE [chen2025grpo] introduces consistency-aware group normalization to mitigate reward variance, while Q-Ponder [cai2025q] adds a pondering controller. From a data perspective, MoDoMoDo formulates a multi-domain mixture optimization that predicts reward distributions and selects optimal curricula [liang2025modomodo]; V-Triune further unifies perception and reasoning tasks within a single triple-objective pipeline, empirically validating that rule-based RL scales to diverse visual signals [ma2025one]. Collectively, these studies indicate that (i) verifiable rewards can serve as a low-cost alternative to human feedback, (ii) group-relative objectives offer higher training stability than token-level PPO on heterogeneous visual tasks, and (iii) curriculum or data-mixture scheduling is emerging as a key ingredient for broad generalization.
@@ 3.1.2 Spatial and 3D Perception
2D perception. Perception centric works applies RL to sharpen object detection, segmentation and grounding without engaging in lengthy chain–of–thought reasoning. Omni-R1 [zhong2025omni] introduces a two-system (global–local) GRPO pipeline that verifies predictions via rule-based metrics, yielding notable improvements on region-level benchmarks and emotion-recognition tasks. DIP-R1 [park2025dip] further decomposes perception into step-wise “inspect $\rightarrow\!$ observe $\rightarrow\!$ act” cycles, where each stage receives deterministic IoU or counting-based rewards to boost fine-grained detection. Perception-R1 [yu2025perception] revisits the effect of GRPO on a spectrum of detection and OCR datasets. Complementing these, VisRL [chen2025visrl] frames intention-guided focus selection as an RL sub-policy, eliminating expensive region labels and consistently outperforming supervised strong baselines on visual grounding tasks.
3D perception. Beyond 2-D, several studies leverage RL to align multimodal models with physically consistent 3-D layouts. MetaSpatial [pan2025metaspatial] employs rendered depth/IoU rewards to refine spatial reasoning for AR/VR scene generation, whereas Scene-R1 [yuan2025scene] couples video-grounded snippet selection with a two-stage grounding policy to learn 3-D scene structure without point-level supervision. At molecular scale, BindGPT [zholus2025bindgpt] treats atom placement as sequential actions and uses binding-affinity estimators as verifiable rewards, demonstrating the scalability of perception-focused RL to 3-D biochemical design. Collectively, these approaches underscore a common recipe: (i) formulate detection/segmentation/3-D alignment as Markov decision problems, (ii) craft deterministic spatial rewards (e.g., IoU, depth consistency, binding energy), and (iii) fine-tune pretrained VLM backbones via GRPO/PPO for stable perception enhancement—thereby differentiating themselves from reasoning-oriented RL variants.
@@ 3.1.3 Image Reasoning
Think about Image. Early think about image works for spatial VQA employs view-consistent or transformation-invariant objectives, such as SVQA-R1 [wang2025svqa] and STAR-R1 [li2025star]. VL-GenRM [zhang2025vl] and RACRO [gou2025perceptual] refine preference data or caption rewards to curb hallucinations. Benchmark-oriented efforts such as EasyARC [unsal2025easyarc] offer procedurally generated, fully verifiable tasks that suit outcome-based reinforcement learning. To mitigate shortcut reliance and improve generality, Visionary-R1 [xia2025visionary] enforces image interpretation before reasoning, whereas UniVG-R1 [bai2025univg] unifies referring, captioning, and detection by coupling a grounding corpus with GRPO fine-tuning. Extensions such as EchoInk-R1 [xing2025echoink] further enrich visual reasoning by integrating audio–visual synchrony under GRPO optimization. Meanwhile, curriculum-driven frameworks—WeThink, G1, GThinker, and Observe-R1 progressively increase task complexity or introduce re-thinking cues (e.g., difficulty ladders, multimodal format constraints) to cultivate deeper and more structured reasoning capabilities in MLLMs. These methods show that language only RL with well-designed visual correctness rewards significantly improves model accuracy, robustness, and out-of-distribution performance.
Think with Image. Early think with image works grounds reasoning via discrete region–level operations: GRIT [fan2025grit] interleaves bounding-box tokens with language and trains under GRPO to maximize both answer correctness and box fidelity, while VILASR [wu2025reinforcing] generalizes this idea to multi-view and video settings, enforcing cross-view spatial consistency. Ground-R1 [cao2025ground] and BRPO [chu2025qwen] adopt two-stage pipelines that first highlight evidence regions (via IoU-based or reflection rewards) before verbal reasoning. A complementary thread explores pixel-space or sequence-level manipulation. Visual Planning [xu2025visual] replaces text chains with imagined image roll-outs rewarded by downstream task success; Pixel Reasoner [su2025pixel] augments the action space with crop, erase and paint primitives and balances exploration through curiosity-driven rewards, whereas DeepEyes [zheng2025deepeyes] shows that end-to-end RL can spontaneously induce such visual thinking behaviours. Finally, TACO [kan2025taco] introduces a think–answer consistency objective that resamples long visual–verbal chains until their intermediate edits align with the final answer. Together, these systems demonstrate that explicitly generating or editing visual artefacts during reasoning optimized via GRPO or R1-style outcome RL—yields more faithful, interpretable and robust image understanding than language only approaches.